Given the population 3 5,7,9,11,13





How many samples can be made from the population with sample size of 3?





Calculate the mean of the sampling distribution





Compute the variance o the sampling distribution

1
Expert's answer
2022-05-05T09:49:06-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\\
=\cfrac{6! } {3! \cdot3! }=\cfrac{4\cdot5\cdot6}{2\cdot3}=20."



Population mean:

"\mu=\cfrac{3+5+7+9+11+13}{6}=8."


Population variance:

"\sigma^2=\sum(x_i-\mu)^2\cdot P(x_i),"

"X-\mu=\\
=\begin{Bmatrix}
 3-8,5-8,7-8,9-8,11-8,13-8
\end{Bmatrix}="

"=\begin{Bmatrix}
-5, - 3,-1,1,3,5
\end{Bmatrix},"

"\sigma^2=(-5)^2\cdot \cfrac{1}{6}+(-3)^2\cdot \cfrac{1}{6}+\\
+(-1)^2\cdot \cfrac{1}{6}+1^2\cdot \cfrac{1}{6}+3^2\cdot \cfrac{1}{6}+5^2\cdot \cfrac{1}{6}=11.667."



Mean of the sampling distribution of sample means:

"\mu_{\bar x} =\mu=8."


Variance of the sampling distribution of sample means:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{11.667}{3}=3.889."